# references.bib

### v2

@@ 3.2 Future Effects of the EJRA Policy
One implication of this model would be that the statistical model in (2) is misspecified, since it implies that there is a constant effect on vacancies from the EJRA whereas the queuing model says that there is a dynamic response that takes time to take full effect. Secondly, the queuing model is treated as exact, whereas the key parameters are uncertain and estimated from data. In that respect, the estimates reported in the Data Report are subject to wide variability as is any social science phenomenon you are trying to forecast 40 years hence.
@@
The results are reported in Table 1 of the Data Report. In the first ten years, there are on average 27.7 fewer new appointments. In the decade from 2053-2062 there are 13.3 fewer vacancies; which is apparently the chosen steady state regime. To put this in perspective, consider Table 3.1, which shows the total number of established academic staff at Cambridge over the period 2013-2023. The fourth column shows that in 2023 a total of 33 members of academic staff left office at the end of their EJRA year, ostensibly due to the EJRA itself, rather than to their own personal preferences. However, from 2022 to 2023, the total number of established staff grew from 1642 to 1669, a net addition of 27. In fact, since 2013, the number of established staff has grown by approximately 1% per year. If we extrapolate that trend in establishment size to 2053/2062, the loss in new vacancies would be more than matched by the the number of new positions every year. One final point to bear in mind when looking at the long term steady state, the State pension age is set to increase to 69 in 2046-2048 due to the increase in longevity.
@@ 4 Conclusion
This paper critically evaluates the findings of the Data Report for the EJRA Review Group, identifying crucial methodological flaws and misinterpretations. The Data Report suggests that the Employer Justified Retirement Age (EJRA) at the University of Cambridge increased job creation rates and contributed positively to the academic workforce. Our analysis reveals several key issues, including inappropriate filters applied to the data, inconsistent treatment of variables, and erroneous conclusions drawn from statistical models.
@@
The Data Report applies more stringent filters to Cambridge data compared to other Russell Group universities (RRG). Certain categories of staff are included in the Cambridge analysis but excluded from the RRG analysis, leading to inappropriate comparisons. Relevant data on job creation and retirements are excluded without clear justification, potentially deflating or inflating the observed job creation rates. Furthermore, the job creation rate is defined differently for Cambridge and the RRG, resulting in incommensurate comparisons.
@@
The Data Report also contains inconsistencies in how variables are treated and measured. The time periods analysed for Cambridge were not directly comparable to those for the RRG, with trends and changes post-2011 not uniformly analysed across all institutions. Key variables such as retirement age and job creation rates are defined and calculated differently, failing to account for structural differences between institutions. Additionally, the Data Report does not uniformly adjust for macroeconomic factors and other external influences, leading to an inaccurate representation of EJRA’s effect.
@@
Importantly, the Difference-in-Differences model used in the Data Report does not adequately control for pre-existing trends and differences between Cambridge and the RRG, resulting in unjustifiable findings on EJRA’s impact on job creation. The Data Report’s claim that EJRA led to a 1.6 percentage point increase in job creation does not hold under proper controls and comparisons. The dynamic model used is inconsistent with the static regression approach, assuming a constant effect over time and ignoring potential variable interactions. The Data Report’s results lack robustness checks, confidence intervals, and measures of uncertainty, with statistical significance disappearing when alternative methods (such as the wild bootstrap) were applied, indicating unreliable conclusions.
@@
Our analysis shows that Cambridge University consistently had lower job creation rates for Established Academic Careers (EAC) compared to the RRG average, both before and after EJRA implementation in 2011. Cambridge’s average job creation rate was approximately 3% per year, while the RRG average was about 5%. There was no relative improvement in job creation rates at Cambridge post-EJRA implementation. Every year since 2011, Cambridge has created fewer EAC jobs than the RRG average, contradicting the Data Report’s claim of a positive impact.
@@
We are grateful to Simon Baron-Cohen for helpful comments.
@@ 5 Appendix
The main model that is used for establishing a significant effect is presented in Section 3.1 of the Data Report, the first displayed equation. We can rewrite this equation in more standard notation as
$r_{it}^{jc}=\alpha_{i}+\gamma_{t}+\delta D_{it}+\varepsilon_{it},$ (2)
The variable $D_{it}$ is the treatment dummy, equal to one for Cambridge post-2012 and equal to zero for Cambridge pre-2012, and equal to zero for all the RRG universities from 2007-2008 to 2021-2022. This approach is called a reversal design and it is used since the retirement policy at Cambridge did not change, while it did at the RRG (it was abolished).
The parameter of interest is $\delta$, which captures the effect of EJRA on job creation (which is implicitly constant across time). The statistical null hypothesis is that $\delta=0$, i.e., that the EJRA had no effect on job creation versus the general alternative. Table A.4 column 1 reports the results of this regression separately for EAC and EAR groups.
@@
In this case there is only one observation in $Cam_{t}$. By comparing with 2012, they carefully use the point of maximum difference to pivot around. A more standard approach would be to compare with the pre-policy average $\overline{Cam}_{pre}-\overline{\overline{RRG}}_{pre}$.
The authors’ main estimates in column 1 of Table A.4. indicate that $\delta$ is positive and statistically significant under their calculated standard errors. However, when they apply the wild bootstrap standard errors in Section 3.4, they report that the statistical significance vanishes. The authors acknowledge that the usual parallel trends assumption is violated in this data, as is clear from Figure A1 and A30. This would imply the estimates in the first column of Table A.4 are suspect. To address this issue, they do the following described on page 9. They fit the regression model
@@
This is absurd. To understand why, imagine constructing a linear trend from five observations on any noisy time series and then extrapolating a further ten observations based on that trend. This is extremely bad practice from a time series point of view, and any public body such as the Bank of England could not argue policy on the basis of such evidence. In Figure A3, one can see two RRG universities have rather extreme boosts in job creation followed by a sharp reversal in one case and slower reversal in the other case, all before 2012. So whatever led them to increase job creation and decrease job creation could be repeated at any time after 2012. Logically, if you extrapolated the trend backwards, it appears that for some universities in 2000, the job creation would be negative. Hence the foundation of this pretrend assumption/method is very weak. Mechanically, however, the result of this process is to drag down the RRG values as those fake pretrends did not continue and this gives a boost to the estimated treatment effect. The treatment effect here is to be interpreted relative to a pre-established trend, so the figure of 4.408 in Table A.4 column 2 is not directly comparable with the result from column 1.
@@
where $s_{it}$ is the log of FTE student numbers at university $i$ in year $t$. In this case, the $\beta$ is assumed constant over universities and again estimated using the pre-2011 data. They then calculate $\widetilde{r}_{it}^{jc}=r_{it}^{jc}-\widehat{\beta}s_{it}$ for all $i,t$ and then apply the DID to this residual.
In this specification, it is odd to use the log of student numbers to predict the ratio of new hires. If we assume that the staff-student ratio is targeted to be constant we would have (where $N$ is the number faculty and $S$ is student numbers):
@@
so that $dN/N$ should be equal to $d\log S$, not $\beta\log S.$ Empirically, $s=\log S$ is more related to $\log N$ than $d\log N$. It is also surprising that in this case, they do not allow $\beta$ to vary with university, perhaps because the result would have been different had they done that. The estimates from this equation are not credible either. Interestingly, in column 3 of Table A.4 they show that for this method the EJRA for EAR staff produces a significant increase in vacancies.
The inference methods are “robust standard errors” in the main part and a wild bootstrap method in Section 3.4. The standard errors in Table A.4 effectively assume that $\varepsilon_{it}$ is independent across $i$ and $t,$ that is, after accounting for the “main effects” $\alpha_{i}$ (university specific effect) and $\gamma_{t}$ (time specific effect), the remaining movements are uncorrelated. This is clearly a bad assumption in this case as the universities in the RRG are not separate individuals whose health outcomes do not affect others’ outcomes but directly competing institutions who are subject to the same rules and procedures (apart from EJRA). The pre-trend estimation is not taken into account of in the standard errors. Instead, the authors just supply the residuals to STATA and get the usual standard errors that would arise if there were no pre-estimation. Clearly, estimating two parameters from five observations is pushing the boundaries of credibility. Credible standard errors would be much larger than those presented. Indeed, as acknowledged in Section 3.4 (robustness checks), when the wild bootstrap is used (this takes account of heteroscedasticity but not cross university and cross time correlations), the statistical significance disappears. Likewise, when a synthetic control method is used, that is, finding a linear combination of RRG that best matches Cambridge in the pre treatment period with respect to vacancy rate, they report that the statistical significance disappears.

### v3

@@ 3.2 Future Effects of the EJRA Policy
One implication of this systems model would be that the statistical model in (2) is misspecified, since it implies that there is a constant effect on vacancies from the EJRA (consistent with the new steady state) whereas the systems dynamic adjustment model says that the response to a change takes time to fully play out to the new steady state. The kind of dynamic responses are shown in their Figure A39, which show some hopping about the steady state line in some cases. Secondly, the systems dynamic model is treated as exact, whereas the key parameters are uncertain and estimated from data. In that respect, the estimates reported in the Data Report are subject to wide variability as is any social science phenomenon you are trying to forecast 40 years hence. Thirdly, the systems model makes strong predictions about how vacancies and employment should have responded when mandatory retirement was abolished in the RRG in 2011, but the actual data do not seem to adjust in the simplistic way that such an idealized system predicts. No attempt was made to validate this as a predictive model on the experience of the RRG universities after the abolishment of their mandatory retirement.
@@
The conclusions drawn by the University’s Review Group are reported in Table 1 of the Data Report. In the first ten years, there are on average 27.7 fewer new appointments. In the decade from 2053-2062 there are 13.3 fewer vacancies; which is apparently the chosen steady state regime. To put this in perspective, consider Table 3.1, which shows the total number of established academic staff at Cambridge over the period 2013-2023. The fourth column shows that in 2023 a total of 33 members of academic staff left office at the end of their EJRA year, ostensibly due to the EJRA itself, rather than to their own personal preferences. However, from 2022 to 2023, the total number of established staff grew from 1642 to 1669, a net addition of 27. In fact, since 2013, the number of established staff has grown by approximately 1% per year. If we extrapolate that trend in establishment size to 2053/2062, the loss in new vacancies would be more than matched by the the number of new positions every year. One final point to bear in mind when looking at the long term steady state, the State pension age is set to increase to 69 in 2046-2048 due to the increase in longevity.
@@ 4 Conclusion
This paper critically evaluates the findings of the Data Report for the EJRA Review Group, identifying crucial methodological flaws and misinterpretations. The Data Report suggests that the Employer Justified Retirement Age (EJRA) at the University of Cambridge increased job creation rates and contributed positively to the academic workforce. Our analysis reveals several key issues, including selective filtering applied to the data, inconsistent treatment of variables, and erroneous conclusions drawn from statistical models.
@@
The Data Report uses internal HR data from Cambridge for defining groups, with filters applied to differentiate between EAC Staff and EAR Staff. These filters include various employment contract details and salary points based on internal Cambridge HR data, which are not necessarily applicable or comparable to other UK HEPs. Relevant data on job creation and retirements are excluded without clear justification, potentially deflating or inflating the observed job creation rates. Furthermore, the job creation rate is defined differently for Cambridge and the RRG, resulting in incommensurate comparisons.
@@
The Data Report also contains inconsistencies in how variables are treated and measured. The time periods analysed for Cambridge were not directly comparable to those for the RRG, with trends and changes post-2011 not uniformly analysed across all institutions. Key variables such as retirement age and job creation rates are defined and calculated differently, failing to account for structural differences between institutions. They also fail to account for the difference between vacancies and jobs, and fail to account for the difference between vacancies due to the EJRA and vacancies due to voluntary retirement at that age. Additionally, the Data Report does not uniformly adjust for macroeconomic factors and other external influences, leading to an inaccurate representation of EJRA’s effect.
@@
Importantly, the Difference-in-Differences model used in the Data Report does not adequately control for pre-existing trends and differences between Cambridge and the RRG, resulting in unjustifiable findings on EJRA’s impact on job creation. The Data Report’s claim that EJRA led to a 1.6 percentage point increase in job creation does not hold under proper controls and comparisons. The dynamical systems model used to forecast the effect of a policy change is inconsistent with the static regression approach, assuming a constant effect over time and ignoring potential variable interactions. The Data Report’s results lack of: robustness checks, valid confidence intervals, and measures of uncertainty, with statistical significance disappearing when alternative methods (such as the wild bootstrap) were applied, suggests that their conclusions are unreliable.
@@
The raw data shows that Cambridge University consistently had lower job creation rates for Established Academic Careers (EAC) compared to the RRG average, both before and after EJRA implementation in 2011. Cambridge’s average job creation rate was approximately 3% per year, while the RRG average was about 5%. There was no relative improvement in job creation rates at Cambridge post-EJRA implementation. Every year since 2011, Cambridge has created fewer EAC jobs than the RRG average, contradicting the Data Report’s claim of a positive impact.
@@
Cambridge has a poor record on discrimination, going back in history to the much delayed acceptance of female graduates, and now to continued age discrimination against academics. As documented in this paper, the Data Report which underpins EJRA is flawed and its claims about EJRA’s contribution to intergenerational fairness, career progression, and succession planning are unsupported. The proposal to abolish EJRA for only academic-related staff is discriminatory against academic staff. Abolishing EJRA also for academic staff will allow senior academics to continue contributing substantially to the university’s research income which supports early career researchers. This will foster a fair and thriving academic community at Cambridge.
We are grateful to Simon Baron-Cohen and Matthew Kramer for helpful comments. We are also grateful to an anonymous researcher involved in the HESA data report for answering our questions.
@@ 5 Appendix
The main model that is used for testing whether a significant effect is present is given in Section 3.1 of the Data Report, the first displayed equation. [^8] There are several typos in the equations and we have verified by personal communication that this is what they did. We can rewrite this equation in more standard notation as
$r_{it}^{jc}=\alpha_{i}+\gamma_{t}+\delta D_{it}+\varepsilon_{it}.$ (2)
The variable $D_{it}$ is the treatment dummy, equal to one for Cambridge post-2012 and equal to zero for Cambridge pre-2012, and equal to zero for all the RRG universities from 2007-2008 to 2021-2022. This approach is called a reversal design and it is used since the retirement policy at Cambridge did not change, while it did at the RRG (it was abolished). The key feature of this model is that there is no interaction between university and time except through the specific role of $D_{it}$. This is clearly a very strong assumption that they themselves show is not credible.
The parameter of interest is $\delta$, which captures the effect of EJRA on job creation (which is implicitly assumed constant across time). The statistical null hypothesis is that $\delta=0$, i.e., that the EJRA had no effect on job creation versus the general alternative that it did. Table A.4 column 1 reports the results of this regression separately for EAC and EAR groups.
@@
In this case there is only one observation in $Cam_{t}$. By comparing with 2012, they are using the point of maximum difference to pivot around. A more standard approach would be to compare with the pre-policy average $\overline{Cam}_{pre}-\overline{\overline{RRG}}_{pre}$.
The authors’ main estimates in column 1 of Table A.4. indicate that $\delta$ is positive and statistically significant under their calculated standard errors. However, when they apply the wild bootstrap standard errors in Section 3.4, they report that the statistical significance vanishes. The authors acknowledge that the usual parallel trends assumption is violated in this data, as is clear from Figures A1 and A30. This would imply the estimates in the first column of Table A.4 are biased. To address this issue, they do the following described on page 9. They fit the regression model
@@
This is absurd. To understand why, imagine constructing a linear trend from five observations on any noisy time series and then extrapolating a further ten observations based on that trend. This is extremely bad practice from a time series point of view, and any public body such as the Bank of England could not argue policy on the basis of such evidence. In Figure A3, one can see two RRG universities have rather extreme boosts in job creation followed by a sharp reversal in one case and slower reversal in the other case, all before 2012. So whatever led them to increase job creation and decrease job creation could be repeated at any time after 2012. Logically, if you extrapolated the trend backwards, it appears that for some universities in 2000, the job creation would be negative. Hence the foundation of this pretrend assumption/method is very weak. Mechanically, the result of this method is to drag down the RRG values as those fake pretrends did not continue and this gives a boost to the estimated treatment effect. The treatment effect here is to be interpreted relative to a pre-established trend, so the number 4.408 in Table A.4 column 2 is not directly comparable with the result from column 1. In addition, the standard errors are calculated incorrectly since they ignore the first stage detrending as if this has no effect on the sampling variability - but it has a huge effect on standard errors so these results are misleading and deceptive.
@@
where $s_{it}$ is the log of FTE student numbers at university $i$ in year $t$. In this case, the $\beta$ is assumed constant over universities and again the estimation of $\beta$ is done using the pre-2011 data. They then calculate $\widetilde{r}_{it}^{jc}=r_{it}^{jc}-\widehat{\beta}s_{it}$ for all $i,t$ and then apply the DID to this residual. The same comments apply to this method in that the preliminary regression on student numbers for five years of data is likely to be very noisy and this noise is not properly taken account of in the standard errors.
Also, the specification seems a bit dubious to say the least. Why use only pre EJRA data to estimate the effect of student numbers on vacancies? Why restrict the parameter $\beta$ to be constant across all universities when we know that student staff ratios vary considerably across universities. In fact, we have done some detective work ourselves to address [Holmes24]. We managed to reconstruct the VCR time series for Cambridge and the RRG from their Figure A1 using specialized software and obtained the student numbers by university and academic year directly from HESA. In our reconstruction of their regression we find that the pooled slope is positive, whereas the slope for Cambridge separately is negative and the slope for RRG separately is positive (none is statistically significant though even if you take the OLS standard errors), which is what they work with, apparently. This is why relative to that pre EJRA trend, Cambridge appears to do well compared to RRG. It is also odd to use the log of student numbers to predict the ratio of new hires. If we assume that the staff-student ratio is targeted to be constant we would have (where $N$ is the number faculty and $S$ is student numbers):
@@
so that $dN/N$ should be equal to $d\log S$, not $\beta\log S.$ Empirically, $s=\log S$ is more related to $\log N$ than $d\log N$. The estimates from this equation are therefore not credible either. Interestingly, in column 3 of Table A.4 they show that for this method the EJRA for EAR staff produces a significant increase in vacancies. If these are the results they wish to hang their clogs on then they would have to change their proposal to abolish EJRA for EAR staff.
For inference, they are using normal critical values in the main part, which can be justified if the difference of vacancy ratios is exactly normally distributed or the relevant sample size is large; here, the binding case is the pre-EJRA period, which has only five observations, which is not usually considered enough for a Central Limit Theorem. The inference methods are “robust standard errors” in the main part (although what they are robust to is not spelled out) and a wild bootstrap method in Section 3.4. The standard errors in Table A.4 effectively assume that $\varepsilon_{it}$ is independent across $i$ and $t,$ that is, after accounting for the “main effects” $\alpha_{i}$ (university specific effect) and $\gamma_{t}$ (time specific effect), the remaining movements are uncorrelated. This is clearly a bad assumption in this case as the universities in the RRG are not separate individuals whose health outcomes do not affect others’ outcomes but directly competing institutions who are subject to the same rules and procedures (apart from EJRA). The pre-trend estimation is not taken into account of in the standard errors. Instead, the authors just supply the residuals to STATA and get the usual standard errors that would arise if there were no pre-estimation. Clearly, estimating two parameters from five observations is pushing the boundaries of credibility. Credible standard errors would be much larger than those presented. Indeed, as acknowledged in Section 3.4 (robustness checks), when the wild bootstrap is used (this takes account of heteroscedasticity but not cross university and cross time correlations), the statistical significance disappears. Likewise, when a synthetic control method is used, that is, finding a linear combination of RRG that best matches Cambridge in the pre treatment period with respect to vacancy rate, they report that the statistical significance disappears.